# Diagrammatic Monte Carlo for finite systems at zero temperature

### v1

@@ Abstract
Figure 1: Approximate number of Feynman diagrams contributing to the self-energy for two and three-body forces. Tha values shown in green give the exact number of Feynman diagrams for the first three orders [28] while valuse at higher order are estimated based on theoir factorial scaling. For each order, very few representative diagrams are shown.
@@
The $\Sigma_{p}^{(\infty)}$ term is a tadpole diagram that plays a special role in the DiagMC algorithm. Its expression includes a convergence factor $e^{i\omega\eta}$ that makes the Monte Carlo integration inefficient; however, it can be easily calculated with standard numerical methods. Moreover, DiagMC calculates the self-energy only up to the overall (real) factor $Z_{p}$ of equation (2). To fix this ambiguity and at the same time address the convergence problem of the tadpole diagrams, we approached the problem in a novel way. First, the $\Sigma_{p}^{(\infty)}$ diagram, and any diagram that includes repetitions of it, are simple enough that they can be evaluated exactly with an analytical frequency integration followed by standard integration over the remaining single-particle quantum numbers. Hence, we compute it separately to avoid any Monte Carlo instability. Second, the Markov chain walker of DiagMC can still be implemented to retain all steps over the tadpole diagrams: simply such contributions are ignored in the sum of Eq. (2) to avoid any double counting. This fact gives us the additional freedom needed to determine the overall normalization. We replace the tadpoles diagrams with unphysical ones that have an exactly known analytical solution and compare it with the partial DiagMC resummation over the same terms. We can deduce the overall normalization factor by counting the times the unphysical diagram is visited compared to the total number of samples [43].
@@ Results.
We perform calculations in the so-called $sc0$ framework [51, 52]. This is a standard prescription in nuclear theory that consists in imposing a fully self-consistent calculation for the static component of the self-energy, $\Sigma_{p}^{(\infty)}$, while the dynamic part is expanded with respect to a mean-field reference propagator. Figure 3 reports the components $p=1$ and $p=6$ of the imaginary part of the irreducible self-energy for $g=0.3$, $\eta=0.1$, and $D=10$. The results were obtained with a Hartree-Fock (HF) reference propagator by sampling ladder diagrams up to order eight. The simulated self-energy shows only small deviations from the exact analytic form (3), so that the NLLSM fit is extremely precise.
@@
Our implementation of DiagMC carries three sources of uncertainty. The first is the statistical error that is intrinsic to any Monte Carlo approach and can be controlled to a good extent by additional sampling. Second, we do perform a many-body truncation by restricting our diagram space up to eighth order in PT and to ladder diagrams only. The error involved is negligible because excluded diagram topologies are known to be irrelevant for the Richardson Hamiltonian [27] and the first eight PT orders are substantially converged. The comparison between the NLLSM fit of Eq. (3) and the DiagMC sampling, shown in Fig. 3, demonstrates the quality of our simulations. Finally, we have a dependence on the regulator parameter $\eta$ [43]. In principle, the exact result is recovered in the limit $\eta\to 0$, where it would cause numerical instabilities. In fact, however, a finite value of $\eta$ simply sets the energy resolution of our simulations, and how ’small’ it should be depends on the specific application.
@@
Figure 5: Correlation energy of the Richardson model at half filling. The errors reported in the plot are $1\%$ of the correlation energy. The regulator was set at $\eta=0.1$. Top panel: Dependence of the correlation energy on the coupling constant $g$ in the model space $D=10$. Bottom panel: Correlation energy per particle for different model space dimensions and $g=0.3$.
The correlation energies derived from the Migdal-Galitskii-Koltun (MGK) sum rule [53, 54] are benchmarked against the exact full configuration interaction (FCI) solutions and the ADC(3) many-body truncation values in Fig. 5. The left panel shows the correlation energy dependence on the coupling constant $g$ within the model space of size $D=10$. The right panel shows the correlation energy per particle with changing model space size $D$, at half-filling, and $g=0.3$. We find agreement with the exact energies within the error bars of the Monte Carlo uncertainty, confirming that the present ladder resummation dominates the self-energy expansion. Remarkably, DiagMC with the HF reference propagator outperforms ADC(3). This discrepancy is to be ascribed to the presence of time inversion terms in the ladder series, which is resummed to all orders in ADC(3) but only in a Tamm-Dancoff fashion [27]. DiagMC correctly resums the Feynman ladders, accounting for all possible time orderings.

### v2

@@ Abstract
Figure 1: Approximate number of Feynman diagrams contributing to the self-energy for two and three-body forces. The values shown in green give the exact number of Feynman diagrams for the first three orders [28] while valuse at higher order are estimated based on their factorial scaling. For each order, very few representative diagrams are shown.
@@
The $\Sigma_{p}^{(\infty)}$ term is a tadpole diagram that plays a special role in the DiagMC algorithm. Its expression includes a convergence factor $e^{i\omega\eta}$ that makes the Monte Carlo integration inefficient; however, it can be easily calculated with standard numerical methods. Moreover, DiagMC calculates the self-energy only up to the overall (real) factor $Z_{p}$ of Eq. (2). To fix this ambiguity and at the same time address the convergence problem of the tadpole diagrams, we approached the problem in a novel way. First, the $\Sigma_{p}^{(\infty)}$ diagram, and any diagram that includes repetitions of it, are simple enough that they can be evaluated exactly with an analytical frequency integration followed by standard integration over the remaining single-particle quantum numbers. Hence, we compute it separately to avoid any Monte Carlo instability. Second, the Markov chain walker of DiagMC can still be implemented to retain all steps over the tadpole diagrams: simply such contributions are ignored in the sum of Eq. (2) to avoid any double counting. This fact gives us the additional freedom needed to determine the overall normalization. We replace the tadpoles diagrams with unphysical ones that have an exactly known analytical solution and compare it with the partial DiagMC resummation over the same terms. We can deduce the overall normalization factor by counting how many times the unphysical diagram is visited compared to the total number of samples [43].
@@ Results.
We perform calculations in the so-called $sc0$ framework [51, 52]. This is a standard prescription in nuclear theory that consists in imposing a fully self-consistent calculation for the static component of the self-energy, $\Sigma_{p}^{(\infty)}$, while the dynamic part is expanded with respect to a mean-field reference propagator. Fig. 3 reports the components $p=1$ and $p=6$ of the imaginary part of the irreducible self-energy for $g=0.3$, $\eta=0.1$, and $D=10$. The results were obtained with a Hartree-Fock (HF) reference propagator by sampling ladder diagrams up to order eight. The simulated self-energy shows only small deviations from the exact analytic form, Eq. (3), so that the NLLSM fit is extremely precise.
@@
Our implementation of DiagMC carries three sources of uncertainty. The first is the statistical error that is intrinsic to any Monte Carlo approach and can be controlled to a good extent by additional sampling. Second, we do perform a many-body truncation by restricting our diagram space up to eighth order in PT and to ladder diagrams only. The error involved is negligible because excluded diagram topologies are known to be irrelevant for the Richardson Hamiltonian [27] and the first eight PT orders are substantially converged. The comparison between the NLLSM fit of Eq. (3) and the DiagMC sampling, shown in Fig. 3, demonstrates the quality of our simulations. Finally, we have a dependence on the regulator parameter $\eta$ [43]. In principle, the exact result is recovered in the limit $\eta\to 0$, where it would cause numerical instabilities. In fact, however, a finite value of $\eta$ simply sets the energy resolution of our simulations, and how ‘small’ it should be depends on the specific application.
@@
Figure 5: Correlation energy of the Richardson model at half-filling. The regulator is set at $\eta=0.1$ and the error of $1\%$ on the correlation energy is too small to be visible in the plot. Top panel: Dependence of the correlation energy on the coupling constant $g$ in the model space $D=10$. Bottom panel: Correlation energy per particle for different model space dimensions and $g=0.3$.
The correlation energies derived from the Migdal-Galitskii-Koltun (MGK) sum rule [53, 54] are benchmarked against the exact full configuration interaction (FCI) solutions and the ADC(3) many-body truncation values in Fig. 5. The top panel shows the correlation energy dependence on the coupling constant $g$ within the model space of size $D=10$. The bottom panel shows the correlation energy per particle with changing model space size $D$, at half-filling, and $g=0.3$. We find agreement with the exact energies within the error bars of the Monte Carlo uncertainty, confirming that the present ladder resummation dominates the self-energy expansion. Remarkably, DiagMC with the HF reference propagator outperforms ADC(3). This discrepancy is to be ascribed to the presence of time inversion terms in the ladder series, which is resummed to all orders in ADC(3) but only in a Tamm-Dancoff fashion [27]. DiagMC correctly resums the Feynman ladders, accounting for all possible time orderings.